# Jailbreaking Quantum Computers

### v1

@@ Abstract
This work presented the first thorough exploration of the attacks on the interface between gate-level and pulse-level quantum circuits and pulse-level quantum circuits themselves. Typically, quantum circuits and programs that execute on quantum computers, are defined using gate-level primitives. However, to improve the expressivity of quantum circuits and to allow better optimization, pulse-level circuits are now often used. The attacks presented in this work leverage the inconsistency between the gate-level description of the custom gate, and the actual, low-level pulse implementation of this gate. By manipulating the custom gate specification, this work proposes numerous attacks: qubit plunder, qubit block, qubit reorder, timing mismatch, frequency mismatch, phase mismatch, and waveform mismatch. This work demonstrates these attacks on the real quantum computer and simulator, and shows that most current software development kits are vulnerable to these new types of attacks. In the end, this work proposes a defense framework. The exploration of security and privacy issues of the rising pulse-level quantum circuits provides insight into the future development of secure quantum software development kits and quantum computer systems.
@@ 1 Introduction
Quantum computing is advancing rapidly with more and bigger quantum computers coming online every year: from quantum computers with one or two qubits two decades ago, to computers with 1121 qubits presently available; and projections for quantum computers with 200 qubits capable of running 100 million gates by the current decade’s end [1]. These existing quantum computers are categorized as Noisy Intermediate-Scale Quantum (NISQ) quantum computers [2] since they do not provide error correction [3] and have noisy qubits and operations. Nevertheless, they show promise in applications like optimization, natural sciences, artificial intelligence, finance, etc [4, 5, 6, 7, 8, 9, 10].
@@
Nowadays, quantum computers are easily accessible through cloud-based services such as IBM Quantum [11], Amazon Bracket [12], or Microsoft Azure [13]. As the utilization of quantum computers is on the rise, the importance of securing quantum circuits and programs that execute on the quantum computers, also becomes increasingly evident. Quantum circuits serve as the foundational building blocks for quantum computation, orchestrating the manipulation and processing of quantum bits. These circuits encode quantum algorithms and enable the execution of complex computations, offering the potential to outperform classical computers in specific tasks. Typical quantum circuits are usually built using quantum gates, which are abstract units that define specific operations on qubits.
@@
However, quantum circuits are not restricted to be defined only using quantum gates. The landscape is evolving beyond conventional gate-level circuits. Recent advancements highlight the substantial advantages of deconstructing gate-level circuits into lower layers, specifically pulses responsible for controlling qubits in quantum computers. Defining circuits using low-level control pulses proves advantageous for tasks such as quantum circuit decomposition, compilation, and optimization [14, 15, 16, 17, 18, 19], as well as enhancing the expressivity and efficiency of quantum circuits [20, 21, 22, 23]. Despite the many benefits of specifying quantum circuits using lower-level control pulses, security and privacy have seldom been touched on. One recent research work introduced higher-energy state attacks, exploiting control pulses to excite qubits into higher-energy levels, and then abusing the properties of higher-energy states to circumvent standard operations in quantum computers [24].
@@
To address research gaps in this field, this paper undertakes the first comprehensive exploration of security and privacy aspects concerning quantum circuits built partly or wholly from low-level pulses rather than purely from gates. In particular, we introduce a new class of attacks, which leverage the inconsistency between the definition of the custom quantum gate, and the low-level implementation of the pulses that realize the gate operations. For the attacks, we assume attackers can clandestinely implant or modify the low-level pulse implementation such that it is not revealed to the high-level gate definition. The proposed attacks are particularly dangerous for two reasons: Firstly, most current quantum software development kits are deficient in the interface between gate-level and pulse-level, and pulse configurations specifying pulse-level gates are in the analog form. These make our attacks less understandable and difficult to verify. Secondly, due to the continuously changing physical properties of quantum computers, pulse configuration needs frequent updates according to hardware properties to maintain proper functionality. This in turn means that it is almost impossible to verify pulses by comparing them to some prior, correct pulses, since even correct pulses will have changed with time. The pulse data is updated possibly as often as the quantum computers are calibrated, providing space for adversaries to install attacks. Both factors make the attacks challenging to detect and prevent.
Figure 1: The proposed attacks. The quantum circuit is defined by the quantum software development kit (SDK) and executed on the quantum computer. The vulnerable quantum circuit includes the custom gate, whose implementation is specified by the pulse. The attacker can perform channel attacks and pulse attacks.
@@
Our attacks can primarily manifest themselves in two places in the quantum circuit stack. The first location is the interface between gate-level and pulse-level circuits, referred to as channel attacks. These attacks, encompassing qubit spoilage, qubit reorder, and qubit block, focus more on software development kit aspects. We scrutinized two widely used quantum SDKs: Qiskit and Amazon Braket Python SDK, and identified the possibility of exploiting channel attacks in both SDKs. The second location is in the pulse configuration, which directly specifies the pulse details. We categorize this as pulse attacks. These attacks include timing mismatch, frequency mismatch, phase mismatch, and waveform mismatch, each exploiting specific features of pulses to manipulate qubits. We demonstrate these attacks on the real quantum computer and simulator, focusing on extensively utilized quantum algorithms: quantum teleportation, Grover’s search, and quantum neural networks.
@@ 2.2 Quantum Gates
A quantum gate $U$ must be unitary, i.e., $UU^{\dagger}=U^{\dagger}U=I$, where $U^{\dagger}$ signifies the conjugate transpose of $U$, and $I$ is the identity matrix. Operating on a qubit |$\ket{\psi}$, a quantum gate $U$ transforms it as |$\ket{\psi}\rightarrow U\ket{\psi}$. Employing the matrix representation, $n$-qubit quantum gates are expressed as $2^{n}\times 2^{n}$ matrices. Quantum gates can be classified as single-qubit gates or multi-qubit gates. U3 gate is one general single-qubit gate that is parameterized by 3 Euler angles. Some other single-qubit gates include Pauli-$X$ gate: a single-qubit gate similar to the classical NOT gate, flips |$\ket{0}$ to |$\ket{1}$ and vice versa; RZ gate: a phase shift between |$\ket{0}$ and |$\ket{1}$. Unlike classical computing, multi-qubit gates can have the ability to create quantum entanglement. One notable example is the CNOT gate, also known as the CX gate, a two-qubit gate that applies a Pauli-$X$ gate to the target qubit if the control qubit is in state |$\ket{1}$, otherwise, it remains unchanged. Another example is the control U3 gate, which also uses one bit to control the application of the U3 gate.
@@ 2.3 Control Pulses
Pulses for all native gates on current quantum computers are often predefined. Due to the constantly changing environment, the physical properties of quantum computers are also continuously varying. As a result, their parameters are regularly updated via automated measurements and calibrations to maintain high fidelity over time.
2.4 Pulse-Level Circuit Description
Quantum circuits described by quantum gates are gate-level circuits. The underlying implementations of the operations on hardware, typically microwave pulses introduced before, are hardware- and technology-dependent. To realize quantum circuits, precise control pulses corresponding to each gate must be generated and transmitted to the qubit. Besides, other information on how they are arranged and where to apply also needs to be specified, such as the starting time steps of pulses, and to which channels pulses should be applied. This comprehensive set of data, or the pulse configuration, along with other pertinent details, constitutes the pulse-level circuit description.
Pulse-level circuits hold significant value for users. First, they facilitate circuit verification and provide insights into execution specifics, revealing the actual operations of quantum circuits. Moreover, pulse-level circuits are actively investigated for optimization and applications as mentioned in Section 1, improving both the expressivity and efficiency.
@@ 2.5 From Gate-Level to Pulse-Level
Typically, quantum circuits and programs are defined using gate-level primitives, with quantum software development kits (SDKs) such as Qiskit [25], Amazon Braket Python SDK [12], Q# [13], Cirq [27].
Similar to compilation in classical computing, gate-level quantum circuits often need to be modified so that they satisfy the requirements of specific hardware. This process involves decomposing non-native quantum gates into native gates, optimizing gates by grouping and eliminating redundant gates, mapping logical qubits to physical qubits, routing circuits within the constrained topology, etc.
The next step is to generate the actual operations, usually pulse-level circuits. This process is like assembling in classical computing. Leveraging pre-calibrated data for native gates on individual qubits or qubit pairs, microwave pulse sequences are generated and tailored for specific hardware. Besides native gates, arbitrary pulses can also be included. Consequently, pulse-level circuits can be readily executed on the target quantum computer.
Pulse-level circuits are necessary because they specify the actual operations on quantum computers. While they are not public on all quantum computers due to reasons like complexity, an increase in exposing it has been seen, including IBM, Amazon, OpenPulse, PennyLane, etc. Though the advantages of pulse-level circuits are widely explored, security and privacy aspects have not been researched, which are the focus of this paper.
@@ 3 Threat Model
In this paper, we introduce attacks targeting quantum circuits that incorporate pulse-level controls. Typically, pulse-level controls are integrated into gate-level circuits using user-defined custom gates. Functioning similar to normal quantum gates, custom gates require specifying the number of qubits and the target qubits for application within a quantum circuit. The detailed operations of custom gates are either independently specified as a series of quantum gates or linked to pulse-level controls. Custom gates can be recursively defined to encompass another custom gate. In this section, we describe the threat model and the stealthiness and classification of the attack.
@@ 3.1 Assumptions
The target of the attacks is quantum circuits that contain custom pulse-level controls. We assume that an attacker is able to modify the specification of the control pulses on victims’ quantum circuits. We assume that after the malicious modifications have been performed unknowingly to victims, victims execute the circuit, thus triggering the operation of the modified pulse gates. In this paper, we assume the attackers can perform malicious modifications, such as modifying the local code to directly change it, or they can leverage malicious toolchains such as malicious compilers to secretly change the pulse information, or they can alter public repositories which victims use to download the code, such as through supply chain attacks.
While it is conceivable for attackers to modify other parts of quantum circuits, and this could be combined with our attacks, we assume attackers only alter pulse configurations for quantum circuits. This assumption is rooted in the clandestine nature of the attack. We posit that detecting the installation of the attack becomes more challenging for victims when only pulse configurations are modified, as opposed to modifications on other parts. For instance, gate-level modifications can be easily detected through methods performing exact match checks of previous and current quantum circuits, such as quantum computer antivirus [28, 29] or maybe simply the hash check. Conversely, modifications on pulse configurations remain more secret.
@@ 4 Proposed Attacks
In this section, we present a classification of our attacks. The pulse-level controls are typically integrated into quantum circuits through user-defined custom gates according to most current quantum quantum software development kits. The custom gate, a gate-level operation similar to a standard quantum gate, can encapsulate pulse-level controls within its structure. Consequently, the attacks can be inserted at two primary places: the interface between the gate-level and pulse-level description, which we refer to as the channel attacks, and the detailed configuration of the pulse-level controls, which we refer to as the pulse attacks. Besides the distinction at the definition level, our attacks can also be performed together with themselves or with other attacks, or be abused as the approach to realize other attacks, which we refer to as the compound attacks. The illustration of our new attacks is shown in Figure 2.
@@ 4.1 Channel Attacks
  • Qubit Plunder: Qubit plunder arises directly from the misalignment between the qubits of the quantum gate and the channels of the underlying pulses. In most quantum quantum software development kits, custom gates can be applied to any qubit set without corresponding checks on the channels to which the pulses are applied. For instance, a custom gate may be designated for qubits 0 and 1, while the underlying pulses could be intended for channels corresponding to qubits 2 and 3. Victims of this attack suffer when only gate-level verification is performed, as channels can be arbitrarily selected. Besides, this attack can be performed on any type of channel. With qubit plunder, a single-qubit quantum gate could maliciously affect control or readout channels, leading to unwanted entanglement or measurements. This attack may be detected when the plundered channels are already occupied with pulses, but the pulse overlap error does not clearly point to the root cause of the problem.
@@ 4.2 Pulse Attacks
  • Timing Mismatch: This type of attack introduces synchronization issues by altering the timing of pulses. One type of timing attack is to change the duration of the pulses and thus change the influence of the pulses, which can cause the operation to be different or also cause the decoherence similar to the qubit block. However, due to the duration of the pulse being part of the waveform of the pulses, we classify this type of attack as waveform mismatch described in the following. The precise timing of pulses and delays is crucial in the pulse definition, complicated further by constraints such as sampling time intervals, acquire alignment, pulse alignment, and granularity of the pulse. For instance, the starting time of a pulse must be the minimum common multiple of acquire alignment and pulse alignment, and the duration of a pulse must be multiples of the pulse granularity. Otherwise, the synchronization problem will cause the qubit to oscillate between |$\ket{0}$ and |$\ket{1}$. Attackers can exploit these constraints by introducing slight delays to displace pulses from the allowed increments, thereby affecting results. Fortunately, while earlier versions of Qiskit and IBM machines overlooked timing checks, subsequent updates have addressed this vulnerability. We list here to warn this type of attack when designing the systems.

### v2

# Security Attacks Abusing Pulse-level Quantum Circuits

@@ Abstract
This work presents the first thorough exploration of the attacks on the interface between gate-level and pulse-level quantum circuits and pulse-level quantum circuits themselves. Typically, quantum circuits and programs that execute on quantum computers, are defined using gate-level primitives. However, to improve the expressivity of quantum circuits and to allow better optimization, pulse-level circuits are now often used. The attacks presented in this work leverage the inconsistency between the gate-level description of the custom gate, and the actual, low-level pulse implementation of this gate. By manipulating the custom gate specification, this work proposes numerous attacks: qubit plunder, qubit block, qubit reorder, timing mismatch, frequency mismatch, phase mismatch, and waveform mismatch. This work demonstrates these attacks on the real quantum computer and simulator, and shows that most current software development kits are vulnerable to these new types of attacks. In the end, this work proposes a defense framework. The exploration of security and privacy issues of the rising pulse-level quantum circuits provides insight into the future development of secure quantum software development kits and quantum computer systems.
@@ 1 Introduction
Quantum computing is advancing rapidly with more and bigger quantum computers coming online every year: from quantum computers with one or two qubits two decades ago, to computers with 1121 qubits presently available; and projections for quantum computers with 200-qubit error corrected modularity capable of running 100 million gates by the current decade’s end [1]. These existing quantum computers are categorized as Noisy Intermediate-Scale Quantum (NISQ) devices [2] since they do not provide error correction [3] and have noisy qubits and operations. Nevertheless, they show promise in applications like optimization, natural sciences, artificial intelligence, finance, etc [4, 5, 6, 7, 8, 9, 10].
Today, quantum computers are easily accessible through cloud-based services such as IBM Quantum [11], Amazon Braket [12], or Microsoft Azure [13]. These services are not just for academics, but actual pay-as-you-go cloud services that anybody can use. Many startups and companies do not have their own quantum computers, and they depend on cloud-based quantum computers to run their, often proprietary, quantum circuits. For example, J.P. Morgan Chase leverages cloud-based Quantinuum H-series to develop algorithms for solving linear systems on quantum hardware [14]. As the utilization of quantum computers is on the rise, the importance of securing quantum circuits and programs that execute on the quantum computers, also becomes increasingly evident.
@@
Quantum circuits serve as the foundational building blocks for quantum computation, orchestrating the manipulation and processing of quantum bits. These circuits encode quantum algorithms and enable the execution of complex computations, offering the potential to outperform classical computers in specific tasks. Typical quantum circuits are usually built using quantum gates, which are abstract units that define specific operations on qubits.
@@
However, quantum circuits are not restricted to be defined only using quantum gates. In fact, quantum gates are only one abstract layer, and they need to be compiled into actual operations on qubits by quantum software development kits (SDKs). For example, quantum gates on superconducting quantum computers are realized by electromagnetic pulses. Recent advancements highlight the substantial advantages of deconstructing gate-level circuits into these lower layers. Defining circuits using low-level control pulses proves advantageous for tasks such as quantum circuit decomposition, compilation, and optimization [15, 16, 17, 18, 19, 20], as well as enhancing the expressivity and efficiency of quantum circuits [21, 22, 23, 24]. Despite the many benefits of specifying quantum circuits using lower-level control pulses, security and privacy have seldom been touched on. One recent research work introduced higher-energy state attacks, exploiting control pulses to excite qubits into higher-energy levels, and then abusing the properties of higher-energy states to circumvent standard operations in quantum computers [25].
@@
To address research gaps in this field, this paper undertakes the first comprehensive exploration of security and privacy aspects concerning quantum circuits built partly or wholly from low-level pulses rather than purely from gates. In particular, we introduce a new class of attacks, which leverage the inconsistency between the definition of the custom quantum gate, and the low-level implementation of the pulses that realize the gate operations. For the attacks, we assume attackers can clandestinely implant or modify the low-level pulse implementation such that it is not revealed to the high-level gate definition. The proposed attacks are particularly dangerous for two reasons: Firstly, most current quantum software development kits are deficient in the interface between gate-level and pulse-level, and pulse configurations specifying pulse-level gates are in the analog form. These make our attacks less understandable and difficult to verify. Secondly, pulse configuration is determined by the hardware properties and environment. Due to reasons such as variations in manufacturing and the continuously changing environment within and near the quantum computer, pulse configuration is distinct and unique for a specific quantum hardware and needs frequent updates to maintain proper functionality. This, in turn, means that pulse specifications are individualized and time-restricted. It is almost impossible to verify pulses by setting an overall standard for pulses for all quantum computers, or by comparing them to some prior, correct pulses, since even correct pulses will have changed with time. The pulse data is updated as often as the quantum computers are calibrated, which can be every day, or even every few hours, providing space for adversaries to install attacks. Both factors result in a unique new attack surface in the quantum software supply chain, which is unique and different from classical computing, making the attacks challenging to detect and prevent.
Figure 1: The workflow of a quantum computer, demonstrating where the channel and pulse attacks proposed in this paper can occur. The quantum software development kit (Quantum SDK) is used to develop Quantum Circuits which are then executed on a Quantum Computer. The proposed attacks target the Quantum Circuit specification and can be deployed, among others, through supply chain attacks.
@@
The high-level workflow schematic of our proposed attacks is shown in Figure 1. Our attacks can primarily manifest themselves in two places in the quantum software development stack. The first location is the interface between gate-level and pulse-level circuits, referred to as channel attacks. These attacks, encompassing qubit spoilage, qubit reorder, and qubit block, focus more on software development kit aspects. We scrutinized two widely used quantum SDKs: Qiskit and Amazon Braket Python SDK, and identified the possibility of exploiting channel attacks in both SDKs. The second location is in the pulse configuration, which directly specifies the pulse details. We categorize this as pulse attacks. These attacks include timing mismatch, frequency mismatch, phase mismatch, and waveform mismatch, each exploiting specific features of pulses to manipulate qubits. We demonstrate these attacks on the real quantum computer and simulator, focusing on extensively utilized quantum algorithms: quantum teleportation, Grover’s search, and quantum neural networks.
@@ 2.2 Quantum Gates
A quantum gate $U$ must be unitary, i.e., $UU^{\dagger}=U^{\dagger}U=I$, where $U^{\dagger}$ signifies the conjugate transpose of $U$, and $I$ is the identity matrix. Operating on a qubit |$\ket{\psi}$, a quantum gate $U$ transforms it as |$\ket{\psi}\rightarrow U\ket{\psi}$. Employing the matrix representation, $n$-qubit quantum gates are expressed as $2^{n}\times 2^{n}$ matrices. Quantum gates can be classified as single-qubit gates or multi-qubit gates. U3 gate is one general single-qubit gate that is parameterized by 3 Euler angles. Some other single-qubit gates include Pauli-$X$ gate: a single-qubit gate similar to the classical NOT gate, flips |$\ket{0}$ to |$\ket{1}$ and vice versa; RZ gate: a phase shift between |$\ket{0}$ and |$\ket{1}$. Unlike classical computing, multi-qubit gates can create quantum entanglement. One notable example is the CNOT gate, also known as the CX gate, a two-qubit gate that applies a Pauli-$X$ gate to the target qubit if the control qubit is in state |$\ket{1}$, otherwise, it remains unchanged. Another example is the control U3 gate, which also uses one bit to control the application of the U3 gate.
@@ 2.3 Control Pulses
Pulses for all native gates on current quantum computers are often predefined, but their parameters are updated periodically. The parameters of pulses are determined by many factors such as qubit properties and environment, and thus they are not fixed over time and are unique for a specific quantum computer: Due to the constantly changing environment, the physical properties of quantum computers also continuously vary. As a result, the parameters of pulses are regularly updated over time via automated measurements and calibrations to maintain high fidelity. Because of this, gate-level circuits need to be recompiled before they are executed so that the most recent pulses representing the gates are used. Apart from changing continuously, due to reasons such as variations in manufacture and environment, the parameters of pulses are also distinct for qubits. Parameters of pulses for different qubits on the same quantum computer, qubits on different quantum computers with the same type and model, have different parameters.
2.4 Pulse-Level Quantum Circuit
Quantum circuits described by quantum gates are called gate-level circuits. The gate-level circuits are transformed into pulse-level circuits by the compiler so that they can be executed on the quantum computer, or the user can directly write pulse-level circuits themselves. To realize quantum circuits, the control pulses must be properly arranged, such as the starting time of the pulses, and to which channels pulses should be applied. All of this information constitutes the pulse-level quantum circuit. Pulse-level circuits hold significant value for users. For example, pulse-level circuits are actively investigated for optimization and applications as mentioned in Section 1, improving both the expressivity and efficiency.
How to incorporate pulses into quantum circuits is the design choice of quantum software development kits. One widely used method is to define the custom gates, and specify what effect these custom gates should have on qubits. Functioning similarly to normal quantum gates, custom gates require the specification of the number of qubits and the target qubits for application within a quantum circuit. The detailed operations of custom gates are either independently specified as a list of quantum gates or linked to specifications of pulse-level controls. Custom gates can be recursively defined to encompass another custom gate.
Pulse-level circuits are necessary because they specify the actual operations on quantum computers. While they are not public on all quantum computers due to complexity, an increase in exposing them has been seen, including IBM Qiskit [28], Amazon Braket [29], PennyLane [30], etc. Though the advantages of pulse-level circuits are widely explored, security and privacy aspects have not been researched, which are the focus of this paper.
@@ 2.5 From Gate-Level to Pulse-Level
Typically, quantum circuits and programs are defined using gate-level primitives, with quantum software development kits such as Qiskit [26], Amazon Braket Python SDK [12], Q# [13], Cirq [31]. Similar to compilation in classical computing, gate-level quantum circuits often need to be modified so that they satisfy the requirements of specific hardware. This process involves decomposing non-native quantum gates into native gates, optimizing gates by grouping and eliminating redundant gates, mapping logical qubits to physical qubits, routing circuits within the constrained topology, etc.
The next step is to transform gate-level circuits to lower-level specifications, usually pulse-level circuits. This process is like assembling in classical computing. Leveraging pre-calibrated data for native gates on individual qubits or qubit pairs, microwave pulse sequences are generated and tailored for specific hardware. Besides native gates, arbitrary pulses can also be included with custom gates. Consequently, pulse-level circuits can be processed by quantum computer controllers to generate the actual operations manipulating qubits on the target quantum computer.
@@ 3 Threat Model
In this paper, we introduce attacks targeting quantum circuits that incorporate pulse-level controls. Typically, pulse-level controls are integrated into gate-level circuits using user-defined custom gates.
@@ 3.1 Assumptions
The target of the attacks is quantum circuits that contain custom pulse-level controls. We assume that attackers are able to modify the specification of the control pulses on victims’ quantum circuits. We assume that after the malicious modifications have been performed unknowingly to victims, victims execute the circuit, thus triggering the operation of the modified pulse gates.
Pulse attacks can be performed in two distinct ways. The first way involves the supply chain attack, where attackers are assumed to perform malicious modifications at various points in the software development process. This could include directly altering local code, tampering with public repositories from which victims download code, or using malicious toolchains, such as compromised compilers, to change pulse information covertly. Importantly, the supply chain attack only requires attackers to be able to modify user-level programs secretly. It does not depend on the ability to alter downstream components in the workflow, nor the ability to interfere with the infrastructure that executes quantum programs. This type of attack will be the main focus of this paper.
The second way involves more powerful attackers who are capable of altering the infrastructure itself. Instead of manipulating user-level quantum programs, these attackers can modify the infrastructure so that the pulse specifications do not produce the expected results. For instance, an attacker might manipulate the environment so that the pulse specifications in the program no longer accurately correspond to the actual conditions in the hardware. This will be left for future research.
While it is conceivable for attackers to modify other parts of quantum circuits, and this could be combined with our attacks, we assume attackers only alter pulse configurations for quantum circuits. This assumption is rooted in the clandestine nature of the attack. We argue that detecting the installation of the attack becomes more challenging for victims when only pulse configurations are modified, as opposed to modifications on other parts. For instance, gate-level modifications can be easily detected through methods performing exact match checks of previous and current quantum circuits, such as quantum computer antivirus [32, 33] or maybe simply the hash check. Conversely, modifications on pulse configurations remain more secret. To the best of our knowledge, no current tool exists that could verify pulse-level quantum circuits, and our work raises the necessity of verifying quantum circuits at the lower level.
@@ 4 Proposed Attacks
In this section, we present a classification of our attacks. The pulse-level controls are typically integrated into quantum circuits through user-defined custom gates according to most current quantum software development kits. The custom gate, a gate-level operation similar to a standard quantum gate, can encapsulate pulse-level controls within its structure. Consequently, the attacks can be inserted at two primary places: the interface between the gate-level and pulse-level description, which we refer to as the channel attacks, and the detailed configuration of the pulse-level controls, which we refer to as the pulse attacks. Besides the distinction at the definition level, our attacks can also be performed together with themselves or with other attacks, or be abused as the approach to realize other attacks, which we refer to as the compound attacks. The illustration of our new attacks is shown in Figure 2.
@@ 4.1 Channel Attacks
  • Qubit Plunder: Qubit plunder arises directly from the misalignment between the qubits of the quantum gate and the channels of the underlying pulses. In most quantum software development kits, custom gates can be applied to any qubit set without corresponding checks on the channels to which the pulses are applied. For instance, a custom gate may be designated for qubits 0 and 1, while the underlying pulses could be intended for channels corresponding to qubits 2 and 3. Victims of this attack suffer when only gate-level verification is performed, as channels can be arbitrarily selected. Besides, this attack can be performed on any type of channel. With qubit plunder, a single-qubit quantum gate could maliciously affect control or readout channels, leading to unwanted entanglement or measurements. This attack may be detected when the plundered channels are already occupied with pulses, but the pulse overlap error does not clearly point to the root cause of the problem.
@@ 4.2 Pulse Attacks
  • Timing Mismatch: This type of attack introduces synchronization issues by altering the timing of pulses. One type of timing attack is to change the duration of the pulses and thus change the influence of the pulses, which can cause the operation to be different or also cause the decoherence similar to the qubit block. However, due to the duration of the pulse being part of the waveform of the pulses, we classify this type of attack as waveform mismatch described in the following. The precise timing of pulses and delays is crucial in the pulse definition, complicated further by constraints such as sampling time intervals, acquire alignment, pulse alignment, and granularity of the pulse. For instance, the starting time of a pulse must be the minimum common multiple of acquire alignment and pulse alignment, and the duration of a pulse must be multiples of the pulse granularity. Otherwise, the synchronization problem will cause the qubit to oscillate between |$\ket{0}$ and |$\ket{1}$. Attackers can exploit these constraints by introducing slight delays to displace pulses from the allowed increments, thereby affecting results. Fortunately, while earlier versions of Qiskit and IBM machines overlooked timing checks, subsequent updates have addressed this vulnerability. We list it here to notify this type of attack when designing the systems.